OpenStudy (anonymous):

Tickets to the movies cost $12 for adults and $8 for children. A group of 14 people went to the movies, and the tickets cost $136. The system of equations models this situation, where x is the number of adults and y is the number of children. How many adults and how many children were in the group? A. 9 adults and 5 children B. 6 adults and 8 children C. 4 adults and 10 children D. 8 adults and 6 children

OpenStudy (anonymous):

the equation to this one is: X+Y=14 12X + 8Y = 136

OpenStudy (anonymous):

hey can someone plz help me i really need it plz i only have one more question after this

OpenStudy (triciaal):

you can divide the 2nd eqn by 4 to work with smaller numbers

OpenStudy (triciaal):

there are at least 5 different methods to solve simultaneous equations. for this one let us us substitution

OpenStudy (anonymous):

i need the answers listen to me i know ur older than me and i cannot have the answers but plz i really need it.

OpenStudy (triciaal):

write the expression for one variable from one equation and substitute this expression for the same variable in the other equation. you will have one equation with one variable. solve for the variable and the value you get put this in the expression to calculate the other value.

OpenStudy (triciaal):

what did you get when you divide the 2nd eqn by 4?

OpenStudy (anonymous):

2

OpenStudy (triciaal):

from eqn 1 let y = 14 - x

OpenStudy (anonymous):

what does that mean

OpenStudy (triciaal):

no when you divide the eqn you divide each term

OpenStudy (anonymous):

what do u mean even boys make clearer sense to me

OpenStudy (triciaal):